# ASE: Practical Acoustic Speed Estimation Beyond Doppler via Sound Diffusion Field

### v1

@@ Abstract
Passive human speed estimation plays a critical role in acoustic sensing. Despite extensive study, existing systems, however, suffer from various limitations: First, previous acoustic speed estimation exploits Doppler Frequency Shifts (DFS) created by moving targets and relies on microphone arrays, making them only capable of sensing the radial speed within a constrained distance. Second, the channel measurement rate proves inadequate to estimate high moving speeds. To overcome these issues, we present ASE, an accurate and robust Acoustic Speed Estimation system on a single commodity microphone. We model the sound propagation from a unique perspective of the acoustic diffusion field, and infer the speed from the acoustic spatial distribution, a completely different way of thinking about speed estimation beyond prior DFS-based approaches. We then propose a novel Orthogonal Time-Delayed Multiplexing (OTDM) scheme for acoustic channel estimation at a high rate that was previously infeasible, making it possible to estimate high speeds. We further develop novel techniques for motion detection and signal enhancement to deliver a robust and practical system. We implement and evaluate ASE through extensive real-world experiments. Our results show that ASE reliably tracks walking speed, independently of target location and direction, with a mean error of 0.13 m/s, a reduction of 2.5x from DFS, and a detection rate of 97.4% for large coverage, e.g., free walking in a 4m $\times$ 4m room. We believe ASE pushes acoustic speed estimation beyond the conventional DFS-based paradigm and will inspire exciting research in acoustic sensing.
@@
Figure 1: ASE vs. DFS. Walking speed $\vec{v}$ can be decomposed to radial speed $v_{r}$ and tangent speed $v_{t}$. DFS can only capture radial speed $v_{r}$, but fails to capture $v_{t}$. Conversely, ASE can capture both $v_{r}$ and $v_{t}$, a complete estimation of $v$.
1 Introduction
@@
Capturing the portraits of indoor human activities is an enduring task in the wide sensing community [29, 79, 71, 89, 46]. Frequently, human subjects are in motion, rendering passive speed estimation one of the most fundamental components in human sensing. At the heart of understanding the physical state of moving subjects, speed provides valuable insights into human behaviors and health. With speed profiles, a broad range of applications can be accommodated, such as gait recognition [80, 13], fall detection [33, 22], human activity recognition [91, 19, 42], tracking [46, 38, 10, 75] and fitness tracking [81, 72, 58], etc.
@@
Particularly, walking speed estimation plays an important role in wellbeing monitoring. It is increasingly perceived as the sixth vital sign [39, 16] which is closely associated with and predictive of one’s health conditions [72]. Slowing walking speed suggests increased frailty, leading to potential physical and cognitive decline [50, 49]. Moreover, walking speed acts as a biomarker for gait recognition [73] and an effective indicator of risky falls [22]. Enabling these applications calls for accurate and robust speed estimation, preferably in a contactless and passive manner.
@@
Passive speed estimation, however, is a long-standing open challenge. Various approaches have been proposed for indoor speed estimation using vision systems [6, 4], wireless signals [86, 69, 72], and acoustic signals [83, 10, 35]. Camera-based approaches, such as VICON motion capture system [6], usually provide the most accurate speed but require complex and expensive hardware setups and are limited within a functional area. Wireless sensing has recently been extensively studied, yet mostly relies on specialized radars [76, 3, 1, 70], or certain WiFi chipsets.
@@
Acoustic devices, especially smart speakers and IoT audio, are now widely and interactively available in our everyday lives, often as plug-and-play devices with co-located microphones and speakers, making acoustic sensing an increasingly hot topic in recent years [89, 14, 15, 53, 61, 60]. The widespread usage of such audio devices hold substantial potential for enabling significant applications, provided that accurate and robust speed estimation can be achieved through acoustic sensing.
@@
Existing acoustic speed estimation methods mostly rely on the Doppler Frequency Shifts (DFS) caused by target movements to derive the speed [84, 33]. These approaches, however, suffer from three fundamental limitations in acoustic speed estimation:
@@
  • The low sound speed imposes an innate limit on the maximum Channel State Information (CSI) rate achievable on an acoustic channel [89], which unfortunately is insufficient for estimating high speeds (e.g., typical walking speed of around 1 m/s).
@@
  • Depending on the specific moving direction and location, Doppler-based approaches can only capture the partial speed projected in a radial direction that creates reflection path length changes and thus frequency shifts [46, 67], as shown in Fig. 1.
@@
  • DFS often only utilizes one or a few reflection paths off the target for sensing, leading to performance degradation with the increase of distance, e.g., over 1 m [28, 89, 29, 68, 35], and thereby confining them for short-range sensing.
Figure 2: OTDM Scheme. (a) Given $s_{1}(t)$, the channel is estimated by frame. (b) Given orthogonal sequences $s_{1}(t)$ and $s_{2}(t)$, we can get two channel estimations at the same time. (c) Orthogonal sequences $s_{1}(t)$ and $s_{2}(t)$, with $s_{2}(t)$ delayed by $\Delta d$, are modulated into one sequence $s(t)$. $H_{1}(t)$ and $H_{2}(t)$ estimated from the two sequences are concatenated into $H(t)$ alternately.
@@
To surpass Doppler-based speed estimation, we build a comprehensive model that can capture the entire speed. Our model builds upon the acoustic diffusion model, as illustrated in Fig. 3. Specifically, we investigate the fundamental properties of sound diffusion indoors and introduce a concise approach for speed estimation. We draw aspiration from previous physical studies of room acoustics [26], which model the diffusive sound propagation. Conceptually, ASE employs the distinct spatial distribution of sound pressure field. The analysis of spatial-temporal properties of the sound field shows that the correlation of the sound energy embodies the speed of a moving target. Although Autocorrelation Function (ACF) has been employed for extracting periodic vital signs [20, 63], building the theoretical and practical bridge between ACF of sound and speed is new and non-trivial. In ASE, we establish the theoretical relationship in the context of the acoustic diffusion field and model the ACF of acoustic CSI as a function of the moving speed. Different from DFS, the proposed model statistically leverages all the reflection signals, and integrates over all possible directions, thus making speed estimation less dependent on the moving location and direction and building a novel foundation for speed estimation with commodity acoustic devices.
We also incorporate several effective designs to transform ASE into a practical system. We employ pseudo-noise code in ASE with nice orthogonality and tolerance to interference for CSI estimation from inaudible sound signals. By careful modulation and filtering, the sensing signals are made hardly audible compared to previous designs like the widely used chirps [61], which are considerably intrusive to human ears in practice. Moreover, we identify robust motion indicators and devise an effective approach that embraces frequency diversity to significantly boost the weak signals for speed estimation, which largely extends the sensing coverage and improves the speed estimation accuracy.
We prototype ASE on commodity audio devices. We conduct comprehensive experiments to evaluate ASE for diverse walking behaviors in real-world indoor scenarios. The results show that ASE achieves a mean speed accuracy of 0.13 m/s with a 90%-tile error of $<0.2$ m/s and an overall detection rate of 97.4% in a 4m $\times$ 4m room, significantly outperforming prior DFS-based approach, which yields a 2.5$\times$ higher mean error under the same conditions. We further conduct case studies on human activity recognition, fall detection, and gait analysis as potential applications by profiling the speeds of different human activities. The superior performance validates ASE and its underlying model as a new paradigm of acoustic speed estimation for many applications.
@@
  • We design and implement ASE system using a single commodity microphone. We incorporate a pipeline of distinct techniques and conduct experiments to validate its superior performance to prior approaches.
@@
2 OTDM Design
We first discuss the contradiction between high speed estimation and acoustic CSI rate. After that, we will present our novel solution of Orthogonal Time Delayed Multiplexing (OTDM) scheme to boost the CSI rate.
2.1 High Speed & Low CSI Rate Dilemma
Initially, we will discuss a universal problem of acoustic speed estimation. To profile speed information, a common practice is to estimate Channel State Information (CSI), which characterizes how sound propagates in a Tx/Rx system. The CSI rate $F_{s}$ refers to how many CSIs we can acquire per second in the system. However, the inadequacy of acoustic CSI rate imposes a fundamental challenge for speed estimation. Conceptually, the lower CSI sampling rates we acquire, the smaller the maximum frequency shift can be observed. Given a CSI rate of $F_{s}$, the maximum frequency shift detectable on the Doppler spectrum is $F_{s}/2$. Assuming a carrier frequency of $f$, the maximum measurable speed is
$v_{\max}=\frac{F_{s}}{2\cdot f}\cdot c,$ (1)
which is only 0.4 m/s considering $f=20$ kHz and $F_{s}=$ 49 Hz. When factoring in noise, this obviously falls short of measuring typical indoor walking speeds.
@@ 2.1 High Speed & Low CSI Rate Dilemma
$F_{s}=\frac{f_{s}}{x/c\cdot f_{s}}=\frac{c}{x}.$ (2)
As we can see, if we want to increase the CSI rate, the only way is to limit the sensing distance. For example, considering the in-air sound speed of 343 m/s and the longest propagation path of 7 meters, the minimum channel measurement interval should be larger than 20 ms, resulting in a CSI rate of 50 Hz. However, even if we narrow this range to 3.5 meters, the maximum possible CSI rate without causing signal mixture only increases to 100 Hz, still insufficient for estimating human walk speed around 1 m/s.
This observation reminds us that current CSI rate cannot support high speed estimation for daily activities (e.g., walking), and it is infeasible to increase CSI rate given the sensing distance for current channel design. To tackle this problem, we present a novel OTDM design in the next section.
Figure 3: Sound Diffusion Model. (a) The sound wave traverses while an object is moving $a\rightarrow b$. (b) The sound wave is diffusive in all directions by irregular reflectors.
2.2 OTDM Transmission Scheme
Based on the above discussion, we find a challenging contradiction for acoustic speed estimation: Given the current channel design, it is impracticable to enhance the CSI rate, while not reducing the sensing range or causing signal mixture. The core of this dilemma is the failure to fully utilize the channel capacity. This situation reminds us of concurrent transmission schemes in network data communication, which are achieved through multiplexing modulations to allow for greater resource utilization. These multiplexing schemes can be roughly divided into Time Division Multiplexing (TDM), Frequency Division Multiplexing (FDM), Code Division Multiplexing (CDM), etc, depending on which dimension is being reused. Notably, Orthogonal Frequency Division Multiplexing (OFDM) is an extension of FDM that divides the frequency band into closely spaced, orthogonal subcarriers. Each subcarrier can be modulated independently, thereby providing a larger data rate within the same frequency bandwidth.
@@ 2.2 OTDM Transmission Scheme
The key idea is, by using orthogonal signals, we can transmit multiple sensing signals concurrently over the same physical channel (i.e., speaker-microphone pair) and perform channel estimation separately, which brings more samples of the channel measurements. If we further shift the multiple orthogonal signals by a certain amount of time, we effectively obtain finer-grained CSI measurements in the time domain, i.e., a higher CSI rate. By OTDM, the duration and interval of each separate sensing signal will not be shortened.
Figure 4: Diffusive Scattering in the room. (a) Illustration of diffusion scattering and specular reflections. The acoustic waves not only experiences specular reflections, but also diffusive scatterings on any non-flat planes. (b) A simulation result of how acoustic wave is reflected and scattered in a common room. The black point emits the acoustic rays with frequency range 17k-20kHz. By employing the rich multi-path semantics, these scatterers are creating ”virtual speakers”, thus creating different views of speed estimation.
@@
Suppose we have two sequences $s_{1}(t)$ and $s_{2}(t)$ with decent orthogonality over time. Then we can transmit them concurrently, while still being able to separate them on the receiver side by using correlation. Previous orthogonal transmission schemes typically transmit synchronized signals. Differently, as shown in Fig. 2, to enhance the CSI rate, we need to delay one sequence by $\Delta d=N_{s}/2$, where $N_{s}$ is the default sensing interval as described in §4.1. Then we can acquire two sets of channel estimation: $H_{1}\in\mathbb{R}^{n}$ and $H_{2}\in\mathbb{R}^{n}$. The two CSI measurements, estimated from the two orthogonal sequences respectively, characterize the physical channel independently, yet at a slightly shifted time. Therefore, by stacking the two series of CSI measurements, i.e., $[H_{1}(1),H_{2}(1),\cdots,H_{1}(n),H_{2}(n)]$, we can double the CSI rate from $1/N_{s}$ to $2/N_{s}$ without shortening the sequence length (i.e., signal duration). Our extensive experiment will show the effectiveness of OTDM design. We will detail the design for the most common case (i.e., one speaker) in §4.2, while our scheme can generalize to more.
3 Acoustic Diffusion Speed Model

### v2

@@ Abstract.
Passive human speed estimation plays a critical role in acoustic sensing. Despite extensive study, existing systems, however, suffer from various limitations: First, the channel measurement rate proves inadequate to estimate high moving speeds. Second, previous acoustic speed estimation exploits Doppler Frequency Shifts (DFS) created by moving targets and relies on microphone arrays, making them only capable of sensing the radial speed within a constrained distance. To overcome these issues, we present ASE, an accurate and robust Acoustic Speed Estimation system on a single commodity microphone. We propose a novel Orthogonal Time-Delayed Multiplexing (OTDM) scheme for acoustic channel estimation at a high rate that was previously infeasible, making it possible to estimate high speeds. We then model the sound propagation from a unique perspective of the acoustic diffusion field, and infer the speed from the acoustic spatial distribution, a completely different way of thinking about speed estimation beyond prior DFS-based approaches. We further develop novel techniques for motion detection and signal enhancement to deliver a robust and practical system. We implement and evaluate ASE through extensive real-world experiments. Our results show that ASE reliably tracks walking speed, independently of target location and direction, with a mean error of 0.13 m/s, a reduction of 2.5x from DFS, and a detection rate of 97.4% for large coverage, e.g., free walking in a 4m $\times$ 4m room. We believe ASE pushes acoustic speed estimation beyond the conventional DFS-based paradigm and inspires exciting research in acoustic sensing. Code is available at https://github.com/aiot-lab/ASE.
copyright: acmlicenseddoi: 10.1145/3749475isbn: 2474-9567/2025/9-ART115journalyear: 2025journal: IMWUTjournalvolume: 9journalnumber: 3article: 115publicationmonth: 9ccs: Human-centered computing Ubiquitous and mobile computing systems and toolsccs: Networks Cyber-physical networks
@@
Figure 1. ASE vs. DFS. Walking speed $\vec{v}$ can be decomposed to radial speed $v_{r}$ and tangent speed $v_{t}$. DFS can only capture radial speed $v_{r}$, but fails to capture $v_{t}$. Conversely, ASE can capture both $v_{r}$ and $v_{t}$, a complete estimation of $v$.
1. Introduction
@@
Capturing the portraits of indoor human activities is an enduring task in the wide sensing community (Li et al., 2022b; Xu et al., 2023; Wu et al., 2022; Zhang et al., 2023b; Qian et al., 2018b). Frequently, human subjects are in motion, rendering passive speed estimation one of the most fundamental components in human sensing. At the heart of understanding the physical state of moving subjects, speed provides valuable insights into human behaviors and health. With speed profiles, a broad range of applications can be accommodated, such as gait recognition (Xu et al., 2019a; Chiang et al., 2020), fall detection (Lian et al., 2021; Hu et al., 2021), human activity recognition (Zheng et al., 2019; Ghosh et al., 2019; Ouyang et al., 2022), tracking (Qian et al., 2018b; Mao et al., 2017; Chen et al., 2017) and fitness tracking (Xu et al., 2019b; Wu et al., 2021; Umair Bin Altaf et al., 2015), etc.
@@
Particularly, walking speed estimation plays an important role in well-being monitoring. It is increasingly perceived as the sixth vital sign (Middleton et al., 2015; Fritz and Lusardi, 2009) which is closely associated with and predictive of one’s health conditions (Wu et al., 2021). Slowing walking speed suggests increased frailty, leading to potential physical and cognitive decline (Rosso et al., 2017; Rasmussen et al., 2019). Moreover, walking speed acts as a biomarker for gait recognition (Wu et al., 2020a) and an effective indicator of risky falls (Hu et al., 2021). Enabling these applications calls for accurate and robust speed estimation, preferably in a contactless and passive manner.
@@
Passive speed estimation, however, is a long-standing open challenge. Various approaches have been proposed for indoor speed estimation using vision systems (Vic, 2023; Mot, 2023), wireless signals (Zhang et al., 2018; Wang et al., 2017; Wu et al., 2021), and acoustic signals (Yun et al., 2015; Chen et al., 2017; Mao et al., 2016). Camera-based approaches, such as the VICON motion capture system (Vic, 2023), usually provide the most accurate speed but require complex and expensive hardware setups and are limited within a functional area. Wireless sensing has recently been extensively studied, yet mostly relies on specialized radars (Xie et al., 2023; IWR, 2023; X4M, 2021; Wang et al., 2020; Yuan et al., 2024, 2025), or certain WiFi chipsets.
@@
Acoustic devices, especially smart speakers and IoT audio, are now widely and interactively available in our everyday lives, often as plug-and-play devices with co-located microphones and speakers, making acoustic sensing an increasingly hot topic in recent years (Zhang et al., 2023b; Fan et al., 2023; Ferlini et al., 2021; Shen et al., 2020; Wang et al., 2019b; Wang and Gollakota, 2019). The widespread usage of such audio devices hold substantial potential for enabling significant applications, provided that accurate and robust speed estimation can be achieved through acoustic sensing.
@@
Existing acoustic speed estimation methods mostly rely on the Doppler Frequency Shifts (DFS) caused by target movements to derive the speed (Yun et al., 2015; Lian et al., 2021). These approaches, however, suffer from three fundamental limitations in acoustic speed estimation:
@@
  • The low sound speed imposes an innate limit on the maximum Channel State Information (CSI) rate achievable on an acoustic channel (Zhang et al., 2023b), which unfortunately, is insufficient for estimating high speeds (e.g., typical walking speed of around 1 m/s).
@@
  • Depending on the specific moving direction and location, Doppler-based approaches can only capture the partial speed projected in a radial direction that creates reflection path length changes and thus frequency shifts (Qian et al., 2018b; Wang et al., 2015), as shown in Fig. 1.
@@
  • DFS often only utilizes one or a few reflection paths off the target for sensing, leading to performance degradation with the increase of distance, e.g., over 1 m (Li et al., 2022a; Zhang et al., 2023b; Li et al., 2022b; Wang et al., 2016; Mao et al., 2016), and thereby confining them for short-range sensing.
@@
To surpass Doppler-based speed estimation, we build a comprehensive model that can capture the entire speed. Our model builds upon the acoustic diffusion model, as illustrated in Fig. 3. Specifically, we investigate the fundamental properties of sound diffusion indoors and introduce a concise approach for speed estimation. We draw aspiration from previous physical studies of room acoustics (Kuttruff, 2000), which model the diffusive sound propagation. Conceptually, ASE employs the distinct spatial distribution of the sound pressure field. The analysis of spatial-temporal properties of the sound field shows that the correlation of the sound energy embodies the speed of a moving target. Although the Autocorrelation Function (ACF) has been employed for extracting periodic vital signs (Gong et al., 2022; Wang et al., 2022), building the theoretical and practical bridge between the ACF of sound and speed is new and non-trivial. In ASE, we establish the theoretical relationship in the context of the acoustic diffusion field and model the ACF of acoustic CSI as a function of the moving speed. Different from DFS, the proposed model statistically leverages all the reflection signals, and integrates over all possible directions, thus making speed estimation less dependent on the moving location and direction and building a novel foundation for speed estimation with commodity acoustic devices.
We also incorporate several effective designs to transform ASE into a practical system. We employ pseudo-noise code in ASE with nice orthogonality and tolerance to interference for CSI estimation from inaudible sound signals. By careful modulation and filtering, the sensing signals are made hardly audible compared to previous designs like the widely used chirps (Wang et al., 2019b), which are considerably intrusive to human ears in practice. Moreover, we identify robust motion indicators and devise an effective approach that embraces frequency diversity to significantly boost the weak signals for speed estimation, which largely extends the sensing coverage and improves the speed estimation accuracy.
[…]
@@
  • We design and implement ASE system using a single commodity microphone. We incorporate a pipeline of distinct techniques and conduct experiments to validate its superior performance over prior approaches.
@@
2. OTDM Design
We first discuss the contradiction between high speed estimation and the acoustic CSI rate. After that, we will present our novel solution of Orthogonal Time Delayed Multiplexing (OTDM) scheme to boost the CSI rate.
2.1. High Speed & Low CSI Rate Dilemma
Initially, we will discuss a universal problem of acoustic speed estimation. To profile speed information, a common practice is to estimate Channel State Information (CSI), which characterizes how sound propagates in a Tx/Rx system. The CSI rate $F_{s}$ refers to how many CSIs we can acquire per second in the system. However, the inadequacy of the acoustic CSI rate imposes a fundamental challenge for speed estimation. Conceptually, the lower CSI sampling rates we acquire, the smaller the maximum frequency shift can be observed. Given a CSI rate of $F_{s}$, the maximum frequency shift detectable on the Doppler spectrum is $F_{s}/2$. Assuming a carrier frequency of $f$, the maximum measurable speed is
(1) $v_{\max}=\frac{F_{s}}{2\cdot f}\cdot c,$
which is only 0.4 m/s considering $f=20$ kHz and $F_{s}=$ 49 Hz. Here $c$ denotes the speed of the sound.When factoring in noise, this obviously falls short of measuring typical indoor walking speeds.
@@ 2.1. High Speed & Low CSI Rate Dilemma
(2) $F_{s}=\frac{f_{s}}{x/c\cdot f_{s}}=\frac{c}{x}.$
As we can see, if we want to increase the CSI rate, the only way is to limit the sensing distance. For example, considering the in-air sound speed of 343 m/s and the longest propagation path of 7 meters, the minimum channel measurement interval should be larger than 20 ms , resulting in a CSI rate of 49 Hz. However, even if we narrow this range to 3.5 meters, the maximum possible CSI rate without causing signal mixture only increases to 100 Hz, still insufficient for estimating human walk speed around 1 m/s.
This observation reminds us that the current CSI rate cannot support high speed estimation for daily activities (e.g., walking), and it is infeasible to increase the CSI rate given the sensing distance for the current channel design. To tackle this problem, we present a novel OTDM design in the next section.
2.2. OTDM Transmission Scheme
Based on the above discussion, we find a challenging contradiction for acoustic speed estimation: Given the current channel design, it is impractical to enhance the CSI rate, while not reducing the sensing range or causing signal mixture. The core of this dilemma is the failure to fully utilize the channel capacity. This situation reminds us of concurrent transmission schemes in network data communication, which are achieved through multiplexing modulations to allow for greater resource utilization. These multiplexing schemes can be roughly divided into Time Division Multiplexing (TDM), Frequency Division Multiplexing (FDM), Code Division Multiplexing (CDM), etc, depending on which dimension is being reused. Notably, Orthogonal Frequency Division Multiplexing (OFDM) is an extension of FDM that divides the frequency band into closely spaced, orthogonal subcarriers. Each subcarrier can be modulated independently, thereby providing a larger data rate within the same frequency bandwidth.
Figure 2. OTDM Scheme. (a) Given $s_{1}(t)$, the channel is estimated by frame. (b) Given orthogonal sequences $s_{1}(t)$ and $s_{2}(t)$, we can get two channel estimations at the same time. (c) Orthogonal sequences $s_{1}(t)$ and $s_{2}(t)$, with $s_{2}(t)$ delayed by $\Delta d$, are modulated into one sequence $s(t)$. $H_{1}(t)$ and $H_{2}(t)$ estimated from the two sequences are concatenated into $H(t)$ alternately.
@@ 2.2. OTDM Transmission Scheme
The key idea is that, by using orthogonal signals, we can transmit multiple sensing signals concurrently over the same physical channel (i.e., speaker-microphone pair) and perform channel estimation separately, which brings more samples of the channel measurements. If we further shift the multiple orthogonal signals by a certain amount of time, we effectively obtain finer-grained CSI measurements in the time domain, i.e., a higher CSI rate. By OTDM, the duration and interval of each separate sensing signal will not be shortened.
Figure 3. Sound Diffusion Model. (a) The sound wave traverses while an object is moving $a\rightarrow b$. (b) The sound wave is diffusive in all directions by irregular reflectors.
@@
Suppose we have two sequences $s_{1}(t)$ and $s_{2}(t)$ with decent orthogonality over time. Then we can transmit them concurrently, while still being able to separate them on the receiver side by using correlation. Previous orthogonal transmission schemes typically transmit synchronized signals. Differently, as shown in Fig. 2, to enhance the CSI rate, we need to delay one sequence by $\Delta d=N_{s}/2$, where $N_{s}$ is the default sensing interval as described in §4.1. Then we can acquire two sets of channel estimation: $H_{1}\in\mathbb{R}^{n}$ and $H_{2}\in\mathbb{R}^{n}$. The two CSI measurements, estimated from the two orthogonal sequences respectively, characterize the physical channel independently, yet at slightly shifted times. Therefore, by stacking the two series of CSI measurements, i.e., $H_{1}(1),H_{2}(1),\cdots,H_{1}(n),H_{2}(n)]$, we can double the CSI rate from $1/N_{s}$ to $2/N_{s}$ without shortening the sequence length (i.e., signal duration). Our extensive experiment will show the effectiveness of OTDM design. We will detail the design for the most common case (i.e., one speaker) in §4.2, while our scheme can generalize to more.
3. Acoustic Diffusion Speed Model